# QOSST: A Highly-Modular Open Source Platform for Experimental Continuous-Variable Quantum Key Distribution

### v1

@@ Abstract
I Introduction
Quantum Key Distribution (QKD) aims at the exchange of a random string of bits – a cryptographic key – between two trusted users, commonly referred to as Alice and Bob, who are linked through an untrusted quantum channel and a public authenticated classical channel. This random bit string can then be used in symmetric encryption ciphers. Crucially, the security of the key exchange is not based on computational assumptions but on the laws of quantum physics leading to information-theoretic security Pirandola et al. (2020).
Two major families of QKD protocols can be distinguished by the degrees of freedom in which the key information is encoded. In Discrete-Variable Quantum Key Distribution (DV-QKD), these are properties of single photons, such as polarisation or phase Bennett and Brassard (1984); Lucamarini et al. (2018), while in Continuous-Variable Quantum Key Distribution (CV-QKD), they are continuous degrees of freedom, such as the quadratures of coherent states Grosshans and Grangier (2002). In terms of practical implementation, an important difference between these two families lies in the detection technique. DV-QKD systems require single-photon detectors, which have limited detection rates due to their dead time and need to be operated in sub-kelvin temperatures to achieve a high efficiency Ip et al. (2008); Eisaman et al. (2011). CV-QKD systems, on the other hand, use coherent optical detectors at room temperature. This reduces the size and cost of the detection apparatus and enables higher bandwidth detectors and hence higher measurement rates. Such detectors are also easier to integrate Liu et al. (2022).
Several experimental demonstrations of CV-QKD have been performed in the past few years, including over long distances Zhang et al. (2020), with high key rates Pan et al. (2022); Roumestan et al. (2024) and with integrated transmitters or receivers Aldama et al. (2023); Piétri et al. (2023); Hajomer et al. (2023). For such protocols that are inspired by classical communication systems, an important part of the implementation complexity comes down to the Digital Signal Processing (DSP) techniques required to recover the symbols at Bob’s side and to correct the impairments coming from the transmission at a low Signal-to-Noise ratio (SNR). Indeed, to limit the amount of information leaked to the eavesdropper, Eve, CV-QKD systems operate in the regime of a few photons per symbol, which is the key difference with classical coherent systems that operate at high SNR. This is common to the several variations of CV-QKD protocols that exist, in terms of symbol shaping, generating a Local Oscillator (LO) for the coherent detection, or sending classical information for frequency and phase compensation.
@@ I Introduction
The rest of this paper is structured as follows: in section II, we describe the protocol and the chosen setup, while in section III, we present our open source software for CV-QKD, QOSST, along with an in-depth explanation of the DSP algorithm. In section IV, an experimental benchmark of the software using our optical setup is presented, before discussing the results and drawing conclusions in sections V and VI respectively.
II Protocol and setup
II.1 Description of the protocol
@@
Security proofs for Gaussian modulation are well established, including also finite-size effects. Despite the complexity in proving security for discrete constellations, a security proof is now also available for any such constellation at the asymptotic limit Denys et al. (2021), and was used in a recent implementation with PCS-QAM modulation Roumestan et al. (2024).
In our work here, we use as baseline protocol the Gaussian Modulated Coherent States (GMCS) GG02 protocol Grosshans and Grangier (2002). After the generation of the symbols on Alice’s side with using a random number generator (ideally a true random number generator such as a Quantum Random Number Generator (QRNG)), we prepare them for transmission applying techniques inspired by classical communications. In particular, the quantum symbols are filtered using a Root-Raised Cosine (RRC) filter to occupy a finite bandwidth $B=R_{s}\cdot(1+\beta_{\mathrm{RRC}})$ where $R_{s}$ is the symbol rate and $\beta_{\mathrm{RRC}}$ is the so-called roll-off factor (taking values between 0 and 1). The RRC filter is also applied on Bob’s side, giving, at the end, a Raised Cosine (RC) filter, which fulfills the Nyquist criteria to minimise inter-symbol interference Proakis (2007).
In addition, the DSP algorithm involves the use of classical signals to perform frame synchronisation, clock recovery, carrier frequency estimation and phase compensation. For the synchronisation, we use a Constant Amplitude Zero AutoCorrelation (CAZAC) sequence, namely a Zadoff-Chu sequence Heimiller (1961); Chu (1972); Zepernick and Finger (2005). This sequence has good correlation properties that allow for an easy synchronisation and is added before the quantum data. For clock recovery, carrier frequency estimation and phase compensation, we use two frequency multiplexed pilot tones, which are complex exponentials at defined frequencies. More details on the RC filter and Zadoff-Chu sequence can be found in appendix E.
@@
After parameter estimation, according to the asymptotic security proof for Gaussian modulation and for reverse reconciliation where Bob’s data is used to form the key, it is possible to bound the information that Eve has acquired using the Holevo bound $\chi_{BE}$ and compute the Secret Key Rate (SKR) using the Devetak-Winter formula Devetak and Winter (2005):
@@
To take into account the finite number of samples used to estimate the values of $T$ and $\xi$ in practice, we use worst-case estimators as follows: assuming that $\xi_{B}$ and $T$ are sampled from a Gaussian probability distribution, we compute the values $x_{B}^{w}$ and $T^{w}$ such that the probability that the actual value of $\xi_{B}$ (resp. $T$) is above $\xi_{B}^{w}$ (resp. below $T^{w}$) is $\varepsilon$, which in our case is chosen to be $\varepsilon=10^{-10}$. We also perform finite-size analysis following the analysis of Leverrier et al. (2010) (see section V for more details).
@@
Finally, in the complete protocol Alice and Bob perform error correction and privacy amplification. Error correction can be done for instance using techniques described in Wang et al. (2018); Gümüş et al. (2021). Privacy amplification compresses the key so that Eve does not hold any meaningful information on it, and is typically done using Toeplitz matrices Bai et al. (2022). These steps are not implemented in QOSST yet and are beyond the scope of this paper.
II.2 Experimental setup
The experimental setup is shown in Fig. 1 together with the protocol steps. The setup for generating the quantum states at Alice’s side is composed of a continuous wave laser (NKT Koheras Basik), an IQ modulator (Exail MXIQER-LN-30), a Modulator Bias Controller (MBC) (Exail MBC-IQ-LAB), an electronic Variable Optical Attenuator (VOA) (Thorlabs V1550PA), followed by a 95:5 beam splitter, with the 95% output going into the monitoring photodiode (Thorlabs PM101A) and the 5% output going into a 90:10 beam splitter, whose 10% output is sent to the output of Alice. Alice is also composed of a control PC and a Digital-to-Analog Converter (DAC) (Teledyne SDR14Tx). The rate of the DAC is $2\;$\displaystyle\mathrm{GSa}\text{/}\mathrm{s}$$ with a bandwidth of $1\;$\displaystyle\mathrm{GHz}\text{/}$$, and it has a 14-bit resolution. The DAC is connected to the control PC using a PCIe connector.
@@ II.2 Experimental setup
The entrance of the setup at Bob’s side is controlled by an optical switch (Thorlabs OSW12-1310E), which can be used to disconnect the quantum channel from the detection system during the calibration, in order to prevent an eavesdropper from tampering with this procedure. This also ensures that noise coming from Alice’s setup is not inadvertently included in the calibration. The output of the switch goes into a 50:50 beam splitter, where the signal is mixed with a continuous wave laser (NKT Koheras Basik) acting as a Local Local Oscillator (LLO). The two outputs of the 50:50 beam splitter go into a Balanced Homodyne Detector (BHD) (Thorlabs PDB480C-AC), whose output is filtered using an analog low-pass filter with a cut-off frequency of $700\;$\displaystyle\mathrm{MHz}\text{/}$$ before being sent to Bob’s control PC through an Analog-to-Digital Converter (ADC) (Teledyne ADQ32) working at $2.5\;$\displaystyle\mathrm{GSa}\text{/}\mathrm{s}$$ and with a bandwidth of $1.5\;$\displaystyle\mathrm{GHz}\text{/}$$. The ADC is connected to the PC using a PCIe connector. Bob’s setup also requires calibration, in particular for the efficiency and the electronic noise (see appendix B for details).
Both the transmitter setup of Alice and the receiver setup of Bob use Polarisation Maintaining (PM) components. The quantum channel between Alice and Bob in our experiments is implemented either with a PM electronic VOA (Thorlabs V1550PA) or with a single mode fiber spool with a length of $25.2\;$\displaystyle\mathrm{km}\text{/}$$ (linear losses: $4.75\;$\displaystyle\mathrm{dB}\text{/}$$, attenuation coefficient: $0.188\;$\displaystyle\mathrm{dB}\text{/}\mathrm{km}$$, total losses with connector: $5.22\;$\displaystyle\mathrm{dB}\text{/}$$), whose characteristics are similar to the fibers deployed in standard optical fiber networks. When using a single mode fiber spool, a manual polarisation controller placed at Bob’s input, before the optical switch, is used to recover the polarisation after the transmission.
Both control PCs are connected to the same local network to provide the classical channel. Authentication on the classical channel is done using the Post Quantum Algorithm for Digital Signatures Falcon Pierre-Alain Fouque, Jeffrey Hoffstein, Paul Kirchner, Vadim Lyubashevsky, Thomas Pornin, Thomas Prest, Thomas Ricosset, Gregor Seiler, William Whyte, Zhenfei Zhang (2020); Thomas Prest (2020); fal , which is expected to be standardised in the NIST Post Quantum Cryptography standardization process NIST . There is no other link between Alice and Bob.
III Description of the open source platform
@@
III.1 Design of the software
The software is written using the Python language Python Software Foundation , and was conceived to be modular. Importantly, it is in principle not limited to a specific hardware and transmission or detection schemes. It might however require some work to adapt the DSP for operations with time-multiplexed pilots or in the pulsed regime.
@@ III.1 Design of the software
Figure 2: Interconnections of the QOSST packages along with the main functionalities. An arrow from block A to block B means that the block A is a dependance of block B. For instance, the arrow from qosst-core to qosst-alice means that qosst-alice uses qosst-core. The details of the packages are provided in the main text.
III.2 Hardware Abstraction Layer
@@ III.3 Core
The configuration file, which controls the whole protocol, is written in TOML Tom Preston-Werner since the Python ecosystem appears to adopt TOML as its default configuration language Brett Cannon, Nathaniel Smith, Donald Stufft (2016); Taneli Hukkinen, Shantanu Jain (2022) and this language is also easily readable and modifiable by both humans and machines Brett Cannon, Nathaniel Smith, Donald Stufft (2016); Nathaniel J. Smith (2016). The configuration holds more than 100 parameters and part of them needs to be optimized to reach the best values of excess noise and secret key rate. An example configuration file is provided with the software and the explanation of all the parameters can be found in the documentation.
@@
Figure 3: Possible modulation schemes in the QOSST software. From left to right: Gaussian (variance: 3, number of points: 100000), 8-PSK (variance: 1), 256-QAM (variance: 1), 256-PCS-QAM (variance: 35, $\nu=0.01$ Roumestan et al. (2024)), 256-Binomial-QAM (variance: 3). The variance parameter does not always represent the actual variance of the generated sequence, but is a parameter that is proportional to the actual variance.
@@ III.4 Alice
In the following, we describe the DSP algorithm for signal generation using Nyquist filters and frequency multiplexed pilots. The first step is the generation of the random symbols from the chosen distribution  [^1] The symbols can also be loaded from a file.. In QOSST, at the time of writing of this paper, the entropy source is a pseudo random number generator from the numpy library Harris et al. (2020); in practice, entropy should be provided by a QRNG and this can be easily integrated in QOSST. The symbols are then upsampled to match the symbol rate $R_{s}$ and filtered using a RRC cosine filter with roll-off factor $\beta_{\mathrm{RRC}}$. The quantum data is then shifted in frequency to the center frequency $f_{\mathrm{shift}}$. Two pilot tones, of frequencies $f_{\mathrm{pilot},1}$ and $f_{\mathrm{pilot},2}$, are frequency multiplexed with the signal. Their frequencies are chosen so that
@@
Figure 4: Power Spectral Densities of the electronic noise, electronic and shot noise and signal data. In this example, we can see that $f_{\mathrm{beat}}$ is around $240\;$\displaystyle\mathrm{MHz}\text{/}$$, and that at this frequency, there is some “low-frequency” noise coming from the setup of Alice. This does not impact the performance of the CV-QKD protocol since it is filtered out. The quantum signal has a bandwidth of $140\;$\displaystyle\mathrm{MHz}\text{/}$$ and is centered at $f_{\mathrm{beat}}+f_{\mathrm{shift}}$ with $f_{\mathrm{shift}}=100\;$\displaystyle\mathrm{MHz}\text{/}$$. The two pilots are placed at the frequencies $f_{\mathrm{beat}}+f_{\mathrm{pilot},1}$ and $f_{\mathrm{beat}}+f_{\mathrm{pilot},2}$ with $f_{\mathrm{pilot},1}=180\;$\displaystyle\mathrm{MHz}\text{/}$$ and $f_{\mathrm{pilot},2}=200\;$\displaystyle\mathrm{MHz}\text{/}$$. The first pilot is slightly more powerful than the second one to be consistent when recovering the frequencies of the pilots. The slight signal at $f_{\mathrm{beat}}-f_{\mathrm{shift}}$, $f_{\mathrm{beat}}-f_{\mathrm{pilot},1}$ and $f_{\mathrm{beat}}-f_{pilot,2}$ is due to imperfections on the sideband suppression.

### v3

@@ Abstract
1 Introduction
Quantum Key Distribution (QKD) aims at the exchange of a random string of bits – a cryptographic key – between two trusted users, commonly referred to as Alice and Bob, who are linked through an untrusted quantum channel and a public authenticated classical channel. This random bit string can then be used in symmetric encryption ciphers. Crucially, the security of the key exchange is not based on computational assumptions but on the laws of quantum physics leading to information-theoretic security [1].
Two major families of QKD protocols can be distinguished by the degrees of freedom in which the key information is encoded. In Discrete-Variable Quantum Key Distribution (DV-QKD), these are properties of single photons, such as polarisation or phase [2, 3], while in Continuous-Variable Quantum Key Distribution (CV-QKD), they are continuous degrees of freedom, such as the quadratures of coherent states [4]. In terms of practical implementation, an important difference between these two families lies in the detection technique. DV-QKD systems require single-photon detectors, which have limited detection rates due to their dead time and need to be operated in sub-kelvin temperatures to achieve a high efficiency [5, 6]. CV-QKD systems, on the other hand, use coherent optical detectors at room temperature. This reduces the size and cost of the detection apparatus and enables higher bandwidth detectors and hence higher measurement rates. Such detectors are also easier to integrate [7].
Several experimental demonstrations of CV-QKD have been performed in the past few years, including over long distances [8], with high key rates [9, 10] and with integrated transmitters or receivers [11, 12, 13]. For such protocols that are inspired by classical communication systems, an important part of the implementation complexity comes down to the Digital Signal Processing (DSP) techniques required to recover the symbols at Bob’s side and to correct the impairments coming from the transmission at a low Signal-to-Noise ratio (SNR). Indeed, to limit the amount of information leaked to the eavesdropper, Eve, CV-QKD systems operate in the regime of a few photons per symbol, which is the key difference with classical coherent systems that operate at high SNR. This is common to the several variations of CV-QKD protocols that exist, in terms of symbol shaping, generating a Local Oscillator (LO) for the coherent detection, or sending classical information for frequency and phase compensation.
@@ 1 Introduction
The rest of this paper is structured as follows: in section 2, we describe the protocol and the chosen setup, while in section 3, we present our open source software for CV-QKD, QOSST, along with an in-depth explanation of the DSP algorithm. In section 4, an experimental benchmark of the software using our optical setup is presented, before discussing the results and drawing conclusions in sections 5 and 6 respectively.
2 Protocol and setup
2.1 Description of the protocol
@@
Security proofs for Gaussian modulation are well established, including also finite-size effects. Despite the complexity in proving security for discrete constellations, a security proof is now also available for any such constellation at the asymptotic limit [14], and was used in a recent implementation with PCS-QAM modulation [10].
In our work here, we use as baseline protocol the Gaussian Modulated Coherent States (GMCS) GG02 protocol [4]. After the generation of the symbols on Alice’s side with using a random number generator (ideally a true random number generator such as a Quantum Random Number Generator (QRNG)), we prepare them for transmission applying techniques inspired by classical communications. In particular, the quantum symbols are filtered using a Root-Raised Cosine (RRC) filter to occupy a finite bandwidth $B=R_{s}\cdot(1+\beta_{\mathrm{RRC}})$ where $R_{s}$ is the symbol rate and $\beta_{\mathrm{RRC}}$ is the so-called roll-off factor (taking values between 0 and 1). The RRC filter is also applied on Bob’s side, giving, at the end, a Raised Cosine (RC) filter, which fulfills the Nyquist criteria to minimise inter-symbol interference [15].
In addition, the DSP algorithm involves the use of classical signals to perform frame synchronisation, clock recovery, carrier frequency estimation and phase compensation. For the synchronisation, we use a Constant Amplitude Zero AutoCorrelation (CAZAC) sequence, namely a Zadoff-Chu sequence [16, 17, 18]. This sequence has good correlation properties that allow for an easy synchronisation and is added before the quantum data. For clock recovery, carrier frequency estimation and phase compensation, we use two frequency multiplexed pilot tones, which are complex exponentials at defined frequencies. More details on the RC filter and Zadoff-Chu sequence can be found in appendix E.
@@
After parameter estimation, according to the asymptotic security proof for Gaussian modulation and for reverse reconciliation where Bob’s data is used to form the key, it is possible to bound the information that Eve has acquired using the Holevo bound $\chi_{BE}$ and compute the Secret Key Rate (SKR) using the Devetak-Winter formula [19]:
@@
To take into account the finite number of samples used to estimate the values of $T$ and $\xi$ in practice, we use worst-case estimators as follows: assuming that $\xi_{B}$ and $T$ are sampled from a Gaussian probability distribution, we compute the values $x_{B}^{w}$ and $T^{w}$ such that the probability that the actual value of $\xi_{B}$ (resp. $T$) is above $\xi_{B}^{w}$ (resp. below $T^{w}$) is $\varepsilon$, which in our case is chosen to be $\varepsilon=10^{-10}$. We also perform finite-size analysis following the analysis of [20] (see section 5 for more details).
@@
Finally, in the complete protocol Alice and Bob perform error correction and privacy amplification. Error correction can be done for instance using techniques described in [21, 22]. Privacy amplification compresses the key so that Eve does not hold any meaningful information on it, and is typically done using Toeplitz matrices [23]. These steps are not implemented in QOSST yet and are beyond the scope of this paper.
2.2 Experimental setup
The experimental setup is shown in Fig. 1 together with the protocol steps. The setup for generating the quantum states at Alice’s side is composed of a continuous wave laser (NKT Koheras Basik), an IQ modulator (Exail MXIQER-LN-30), a Modulator Bias Controller (MBC) (Exail MBC-IQ-LAB), an electronic Variable Optical Attenuator (VOA) (Thorlabs V1550PA), followed by a 95:5 beam splitter, with the 95% output going into the monitoring photodiode (Thorlabs PM101A) and the 5% output going into a 90:10 beam splitter, whose 10% output is sent to the output of Alice. Alice is also composed of a control PC and a Digital-to-Analog Converter (DAC) (Teledyne SDR14Tx). The rate of the DAC is $2\;$\mathrm{GSa}\text{/}\mathrm{s}$$ with a bandwidth of $1\;$\mathrm{GHz}\text{/}$$, and it has a 14-bit resolution. The DAC is connected to the control PC using a PCIe connector.
@@ 2.2 Experimental setup
The entrance of the setup at Bob’s side is controlled by an optical switch (Thorlabs OSW12-1310E), which can be used to disconnect the quantum channel from the detection system during the calibration, in order to prevent an eavesdropper from tampering with this procedure. This also ensures that noise coming from Alice’s setup is not inadvertently included in the calibration. The output of the switch goes into a 50:50 beam splitter, where the signal is mixed with a continuous wave laser (NKT Koheras Basik) acting as a Local Local Oscillator (LLO). The two outputs of the 50:50 beam splitter go into a Balanced Homodyne Detector (BHD) (Thorlabs PDB480C-AC), whose output is filtered using an analog low-pass filter with a cut-off frequency of $700\;$\mathrm{MHz}\text{/}$$ before being sent to Bob’s control PC through an Analog-to-Digital Converter (ADC) (Teledyne ADQ32) working at $2.5\;$\mathrm{GSa}\text{/}\mathrm{s}$$ and with a bandwidth of $1.5\;$\mathrm{GHz}\text{/}$$. The ADC is connected to the PC using a PCIe connector. Bob’s setup also requires calibration, in particular for the efficiency and the electronic noise (see appendix B for details).
Both the transmitter setup of Alice and the receiver setup of Bob use Polarisation Maintaining (PM) components. The quantum channel between Alice and Bob in our experiments is implemented either with a PM electronic VOA (Thorlabs V1550PA) or with a single mode fiber spool with a length of $25.2\;$\mathrm{km}\text{/}$$ (linear losses: $4.75\;$\mathrm{dB}\text{/}$$, attenuation coefficient: $0.188\;$\mathrm{dB}\text{/}\mathrm{km}$$, total losses with connector: $5.22\;$\mathrm{dB}\text{/}$$), whose characteristics are similar to the fibers deployed in standard optical fiber networks. When using a single mode fiber spool, a manual polarisation controller placed at Bob’s input, before the optical switch, is used to recover the polarisation after the transmission.
Both control PCs are connected to the same local network to provide the classical channel. Authentication on the classical channel is done using the Post Quantum Algorithm for Digital Signatures Falcon [24, 25, 26], which is expected to be standardised in the NIST Post Quantum Cryptography standardization process [27]. There is no other link between Alice and Bob.
3 Description of the open source platform
@@
3.1 Design of the software
The software is written using the Python language [28], and was conceived to be modular. Importantly, it is in principle not limited to a specific hardware and transmission or detection schemes. It might however require some work to adapt the DSP for operations with time-multiplexed pilots or in the pulsed regime.
@@ 3.1 Design of the software
Figure 2: Interconnections of the QOSST packages along with the main functionalities. An arrow from block A to block B means that the block A is a dependence of block B. For instance, the arrow from qosst-core to qosst-alice means that qosst-alice uses qosst-core. The details of the packages are provided in the main text.
3.2 Hardware Abstraction Layer
@@ 3.3 Core
The configuration file, which controls the whole protocol, is written in TOML [29] since the Python ecosystem appears to adopt TOML as its default configuration language [30, 31] and this language is also easily readable and modifiable by both humans and machines [30, 32]. The configuration holds more than 100 parameters and part of them needs to be optimized to reach the best values of excess noise and secret key rate. An example configuration file is provided with the software and the explanation of all the parameters can be found in the documentation.
@@
Figure 3: Possible modulation schemes in the QOSST software. From left to right: Gaussian (variance: 3, number of points: 100000), 8-PSK (variance: 1), 256-QAM (variance: 1), 256-PCS-QAM (variance: 35, $\nu=0.01$ [10]), 256-Binomial-QAM (variance: 3). The variance parameter does not always represent the actual variance of the generated sequence, but is a parameter that is proportional to the actual variance.
@@ 3.4 Alice
In the following, we describe the DSP algorithm for signal generation using Nyquist filters and frequency multiplexed pilots. The first step is the generation of the random symbols from the chosen distribution  [^1] The symbols can also be loaded from a file.. In QOSST, at the time of writing of this paper, the entropy source is a pseudo random number generator from the numpy library [33]; in practice, entropy should be provided by a QRNG and this can be easily integrated in QOSST. The symbols are then upsampled to match the symbol rate $R_{s}$ and filtered using a RRC cosine filter with roll-off factor $\beta_{\mathrm{RRC}}$. The quantum data is then shifted in frequency to the center frequency $f_{\mathrm{shift}}$. Two pilot tones, of frequencies $f_{\mathrm{pilot},1}$ and $f_{\mathrm{pilot},2}$, are frequency multiplexed with the signal. Their frequencies are chosen so that
@@
Figure 4: Power Spectral Densities of the electronic noise, electronic and shot noise and signal data. In this example, we can see that $f_{\mathrm{beat}}$ is around $240\;$\mathrm{MHz}\text{/}$$, and that at this frequency, there is some “low-frequency” noise coming from the setup of Alice. This does not impact the performance of the CV-QKD protocol since it is filtered out. The quantum signal has a bandwidth of $140\;$\mathrm{MHz}\text{/}$$ and is centered at $f_{\mathrm{beat}}+f_{\mathrm{shift}}$ with $f_{\mathrm{shift}}=100\;$\mathrm{MHz}\text{/}$$. The two pilots are placed at the frequencies $f_{\mathrm{beat}}+f_{\mathrm{pilot},1}$ and $f_{\mathrm{beat}}+f_{\mathrm{pilot},2}$ with $f_{\mathrm{pilot},1}=180\;$\mathrm{MHz}\text{/}$$ and $f_{\mathrm{pilot},2}=200\;$\mathrm{MHz}\text{/}$$. The first pilot is slightly more powerful than the second one to be consistent when recovering the frequencies of the pilots. The slight signal at $f_{\mathrm{beat}}-f_{\mathrm{shift}}$, $f_{\mathrm{beat}}-f_{\mathrm{pilot},1}$ and $f_{\mathrm{beat}}-f_{pilot,2}$ is due to imperfections on the sideband suppression.